A firm maintains a capital structure with debt-equity ratio of 2/3 and has $24,000 earnings for the year-end. The firm follows a residual dividend policy.

a. What is the maximum capital budget amount the firm can finance without raising external equity?

b. Suppose the firm's capital structure before financing the maximum capital budget consisted of $500,000 debt and $750,000 equity. Calculate the firm's debt ratio after financing the maximum capital budget in part (a).
Answer #1

Answer (a):

Given that:

Debt equity Ratio = 2:3

Earnings = $24,000

The firm follows a residual dividend policy.

Debt financing of the capital budget = 24000 * 2/3 = $16,000

Maximum capital budget amount the firm can finance without raising external equity = 16000 + 24000 = $40,000

Maximum capital budget amount the firm can finance without raising external equity = $40,000

Answer (b):

Maximum capital budget = $40,000

Financed by:

Retained earnings = $24,000

Debt = $16,000

Assuming the firm's capital structure before financing the maximum capital budget consisted of $500,000 debt and $750,000 equity.

Post financing:

Debt = 500000 + 16000 = $516,000

Equity = 750000 + 24000 = $774,000

Total assets = 516000 + 724000 = $1,290,000

Firm's debt ratio after financing the maximum capital budget in part (a) = Laibilties / Total assets

= 516000 / 1290000

= 0.40 or 40%

Firm's debt ratio after financing the maximum capital budget in part (a) = 0.40 or 40%